# Elliptic curves and Fourier coefficients of meromorphic modular forms

### v1

@@ 5.3. Hypergeometric reinterpretation of $a_{p^{l}}(F_{4,c})$
Now, one can relate $a_{\mathfrak{p}}(C)$ and $a_{\mathfrak{P}}(C_{1})$ via the fact that $j(C)=j(C_{1})$ and use an appropriate version of the truncated Clausen’s formula to obtain the desired congruence (see e.g. [CDL+13, Lemma 18] for the truncated Clausen’s formula over $\mathbb{Q}$).
6. Proof of Theorem 4.7 and the $\frac{k-2}{2}$-magnetic property
We will now prove Theorem 4.7, i.e. the supercongruences satisfied by $G_{k,D}^{(k/2)}$. With extra work for primes $p$ with $p{\hskip 1.0pt}|{\hskip 1.0pt}D$, it also follows that, up to some integer multiple, $G_{k,D}^{(k/2)}$ is $\frac{k-2}{2}$-magnetic. In fact, we will prove a more general result on supercongruences and the $\frac{k-2}{2}$-magnetic property which holds for any discriminant $D<0$.
Recall that for $k\in\{4,6,8,10,14\}$,
$G_{k}(z,\tau)\;:=\;\frac{E_{k}(z)\cdot\frac{E_{14-k}}{\Delta}(\tau)}{j(z)-j(\tau)}.$
Let $D<0$ be a discriminant. Let $\mathcal{Q}_{D}$ be a complete set of $\mathrm{SL}_{2}(\mathbb{Z})$-inequivalent positive definite quadratic forms $Q(X,Y)\in\mathbb{Z}[X,Y]$ of discriminant $D$ and let $\mathcal{Q}_{D}^{\mathrm{prim}}\subseteq\mathcal{Q}_{D}$ be the subset consisting of primitive elements, i.e., $aX^{2}+bXY+cY^{2}$ with $(a,b,c)=1$. For $Q(X,Y)=aX^{2}+bXY+cY^{2}\in\mathcal{Q}_{D}$, define
$\alpha_{Q}\;:=\;\frac{-b+\sqrt{D}}{2a}\in\mathcal{H}.$
Now, define
$\mathcal{G}_{k,D}(z)\;:=\;\frac{2}{w_{D}}\sum_{Q\in\mathcal{Q}_{D}^{\mathrm{prim}}}(\partial^{(k-2)/2}_{\tau}G_{k})(z,\alpha_{Q}),$
where
$w_{D}\;:=\;\begin{cases}4&\text{ if $D=-4$,}\\ 6&\text{ if $D=-3$,}\\ 2&\text{ otherwise.}\end{cases}$
Note that $(\partial^{(k-2)/2}_{\tau}G_{k})(z,\tau)$ is of weight 0 with respect to $\tau$, so $\mathcal{G}_{k,D}(z)$ is independent of the choice of $\mathcal{Q}_{D}$ (unlike $G_{k,D}^{(r)}$).
@@
Let $k\in\{4,6,8,10,14\}$ and let $D<0$ be a discriminant. Then, $\mathcal{G}_{k,D}$ satisfies the following.
@@
    $\mathcal{G}_{k,D}$ has a Fourier series expansion of the form
    $\mathcal{G}_{k,D}(z)\;=\;\sum_{n=1}^{\infty}a_{n}(\mathcal{G}_{k,D})q^{n}\in\mathbb{Q}[[q]].$
  2. (2)
    Write $D=A^{2}D_{0}$ for some fundamental discriminant $D_{0}<0$ and let
[…]
    Then, $\widetilde{\mathcal{G}}_{k,D}$ is $\frac{k-2}{2}$-magnetic, i.e., $\widetilde{\mathcal{G}}_{k,D}\in\mathbb{Z}[[q]]$ and for all $n\in\mathbb{Z}^{+}$,
@@
    $a_{np^{l}}(\widetilde{\mathcal{G}}_{k,D})\;\equiv\;\big(\big(\tfrac{D}{p}\big)p\big)^{\frac{k-2}{2}}a_{np^{l-1}}(\widetilde{\mathcal{G}}_{k,D})\pmod{p^{(k-1)l}}$.
Theorem 6.1 should be viewed as a natural generalization of Li–Neururer [LN19, Theorem 1.5] and Paşol–Zudilin [PZ22, Theorem 1 and 2]. Indeed, the proof uses the same method: one considers the half-integral weight preimage of $\mathcal{G}_{k,D}$ under the Shimura lift, analyzes the $U_{p}$ action on the preimage, and transfers the properties of the preimage to the supercongruences and the magnetic property of $\mathcal{G}_{k,D}$.
@@ 6.1. The Shimura lift
We first recall some definitions and notations regarding the Shimura lift. For $k,s\in\mathbb{Z}$, let $M_{k}^{!}$ denote the space of weakly holomorphic modular forms of weight $k$ and level $1$ and let $M_{s+1/2}^{!,+}$ denote the Kohnen plus space, i.e., the space of weakly holomorphic modular forms of weight $s+1/2$ and level $\Gamma_{0}(4)$ whose $n$-th Fourier coefficient is zero if $(-1)^{s}n\not\equiv 0,1\pmod{4}$. For $f\in M_{s+1/2}^{!,+}$ and a fundamental discriminant $d_{0}$ with $(-1)^{s}d_{0}>0$, define the $d_{0}$-th Shimura lift of $f$ by
@@
Let $d$ be a discriminant and $d_{0}$ be a fundamental discriminant with $dd_{0}<0$. Let $\mathcal{Q}_{dd_{0}}$ be a complete set of $\mathrm{SL}_{2}(\mathbb{Z})$-inequivalent positive definite quadratic forms $Q(X,Y)\in\mathbb{Z}[X,Y]$ of discriminant $dd_{0}$. For each $Q(X,Y)=aX^{2}+bXY+cY^{2}\in\mathcal{Q}_{dd_{0}}$, define
$\alpha_{Q}\;:=\;\frac{-b+\sqrt{dd_{0}}}{2a}\in\mathcal{H},$
define
@@
and let
[…]
For an $\mathrm{SL}_{2}(\mathbb{Z})$-invariant function $f(\tau)$ on $\mathcal{H}$, recall the following definition from [DJ08b]
(14) $\mathrm{Tr}_{d,d_{0}}(f)\;=\;\sum_{Q\in\mathcal{Q}_{dd_{0}}}\frac{1}{w_{Q}}\cdot\chi(Q)f(\alpha_{Q}).$
@@
Let $2s\in\{4,6,8,10,14\}$. Let $d$ be a discriminant and $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$ and $dd_{0}<0$. Then,
$\big(\mathcal{S}_{d_{0}}f_{s+1/2,|d|}\big)(z)\;=\;-(-1)^{\lfloor\frac{s-1}{2}\rfloor}{\hskip 1.0pt}|d|^{-\frac{s}{2}}{\hskip 1.0pt}|d_{0}|^{\frac{s-1}{2}}\;\mathrm{Tr}_{d,d_{0}}\big((\partial^{s-1}_{\tau}G_{2s})(z,\tau)\big),$
@@
This is a special case of [DJ08b, Proposition 10]. ∎
6.2. The $U_{p}$ action
For a $q$-series $f=\sum_{n\gg-\infty}a_{n}(f)q^{n}$, we adopt the notations $f|U_{p},f|V_{p}$, and $f|T_{p,k}$ for from Section 5. For a map $\chi:\mathbb{Z}\rightarrow\mathbb{C}$ (in our case a Dirichlet character), define
$f|\chi\;:=\;\sum_{n\gg-\infty}\chi(n)a_{n}(f)q^{n}.$
Now, let $s\in\mathbb{Z}$ and let
For a prime $p$, define
Indeed, if $p$ is odd, then
$f|T_{p,s+1/2}\;=\;f|(U_{p}^{2}+p^{s-1}\chi_{p}^{\prime}+p^{2s-1}V_{p}^{2}),$
where $\chi_{p}^{\prime}(n)=\big(\frac{(-1)^{s}n}{p}\big)$.
Now, let $2s\in\{4,6,8,10,14\}$. For every $m\in\mathbb{Z}$ with $m\geq 0$ and $(-1)^{s-1}m\equiv 0,1\pmod{4}$, there exists a unique $f_{s+1/2,m}\in M_{s+1/2}^{!,+}$ such that
$f_{s+1/2,m}\;=\;q^{-m}+O(q)\in\mathbb{Z}[[q,q^{-1}]].$
In particular, these $f_{s+1/2,m}$ form a basis of $M_{s+1/2}^{!,+}$ (see [DJ08b, Section 2 and Appendix] for a discussion of this fact).
@@ 6.2. The $U_{p}$ action
Let $2s\in\{4,6,8,10,14\}$, $p$ be a prime, and $r\in\mathbb{Z}^{+}$. Let $f\in M_{s+1/2}^{!,+}$ be $p$-integral and suppose that $p^{r}{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}(f)$ for all $n\leq 0$. Then, $f\equiv 0\pmod{p^{r}}$.
Proof.
This essentially follows from the argument in the last part of the proof of [PZ22, Lemma 2], which we will recall here. As $(f_{s+1/2,m})_{m}$ forms an integral basis of $M_{s+1/2}^{!,+}$, we must have
$f\;=\;\sum_{n\leq 0}a_{n}(f)f_{s+1/2,-n},$
so $f\equiv 0\pmod{p^{r}}$. ∎
Lemma 6.4.
Let $2s\in\{4,6,8,10,14\}$, $m\in\mathbb{Z}^{+}$ with $(-1)^{s-1}m\equiv 0,1\pmod{4}$, and $p$ be a prime. Suppose that
@@
$g_{0}$ $\;:=\;p^{s-1}\big(\tfrac{(-1)^{s-1}m}{p}\big)\cdot f_{s+1/2,m}$
$g_{1}$ $\;:=\;f_{s+1/2,m}$
$g_{i+1}$ $\;:=\;\frac{1}{p^{2s-1}}\cdot\big(g_{i}|T_{p,s+1/2}-g_{i-1}\big).$
@@
$g_{i}\;=\;q^{-mp^{2i-2}}+O(q)\;=\;f_{s+1/2,mp^{2i-2}}\in\mathbb{Z}[[q]].$
@@
Write $f_{m}=f_{s+1/2,m}$. The case when $i=1$ follows from the definition of $f_{m}$. For $i=2$, it is easy to check that the assumptions on $p$ and $m$ imply that
$g_{1}|T_{p,s+1/2}-g_{0}$ $\;=\;\big(p^{2s-1}q^{-mp^{2}}+p^{s-1}\big(\tfrac{(-1)^{s-1}m}{p}\big)q^{-m}+O(q)\big)-\big(p^{s-1}\big(\tfrac{(-1)^{s-1}m}{p}\big)q^{-m}+O(q)\big)$
$\;=\;p^{2s-1}q^{-mp^{2}}+O(q).$
It then follows from Lemma 6.3 that
$g_{2}\;=\;\frac{1}{p^{2s-1}}\cdot\big(g_{1}|T_{p,s+1/2}-g_{0}\big)\;=\;q^{-mp^{2}}+O(q)\in\mathbb{Z}[[q,q^{-1}]].$
Now, we proceed by induction. For $i\geq 2$, by induction hypothesis,
$g_{i}|T_{p,s+1/2}-g_{i-1}$ $\;=\;\big(p^{2s-1}q^{-mp^{2i}}+q^{-mp^{2i-4}}+O(q)\big)-\big(q^{-mp^{2i-4}}+O(q)\big)$
$\;=\;p^{2s-1}q^{-mp^{2i}}+O(q).$
It then follows from Lemma 6.3 that
$g_{i+1}\;=\;\frac{1}{p^{2s-1}}\cdot\big(g_{i}|T_{p,s+1/2}-g_{i-1}\big)\;=\;q^{-mp^{2i}}+O(q)\in\mathbb{Z}[[q,q^{-1}]].$
Now, recall that the Shimura lift is Hecke-equivariant. That is, for $s\in\mathbb{Z}$, $f\in M_{s+1/2}^{!,+}$, and a prime $p$
$\mathcal{S}_{d_{0}}(f|T_{p^{2},s+1/2})\;=\;\mathcal{S}_{d_{0}}(f)|T_{p,2s},$
@@
Let $2s\in\{4,6,8,10,14\}$ and let $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$. Let $m\in\mathbb{Z}$ with $m\geq 0$ and $(-1)^{s-1}m\equiv 0,1\pmod{4}$ and let $p$ be a prime such that
@@
Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{m})\;\equiv\;p^{s-1}\big(\tfrac{(-1)^{s-1}m}{p}\big)\cdot a_{np^{l-1}}(F_{m})\pmod{p^{(2s-1)l}}$.
@@
$G_{0}$ $\;=\;p^{s-1}\big(\tfrac{(-1)^{s-1}m}{p}\big)\cdot F_{m}$
$G_{1}$ $\;=\;F_{m}$
$G_{i+1}$ $\;=\;\frac{1}{p^{2s-1}}\cdot\big(G_{i}|T_{p,2s}-G_{i-1}\big)\;=\;\frac{1}{p^{2s-1}}\big(G_{i}|U_{p}-G_{i-1}\big)+G_{i}|V_{p}\in\mathbb{Z}[[q]]$
Write $H_{i}=G_{i}|U_{p}-G_{i-1}$. Then,
$H_{i+1}\;=\;G_{i+1}|U_{p}-G_{i}\;=\;\bigg(\frac{1}{p^{2s-1}}\big(G_{i}|U_{p}-G_{i-1}\big)+G_{i}|V_{p}\bigg)\bigg|U_{p}-G_{i}\;=\;\frac{1}{p^{2s-1}}{\hskip 1.0pt}H_{i}|U_{p}.$
By induction, it follows that
$H_{1}|U_{p}^{l}\;=\;p^{(2s-1)l}H_{l+1}.$
@@
$F_{m}|U_{p}^{l}-p^{s-1}\big(\tfrac{(-1)^{s-1}m}{p}\big)\cdot F_{m}|U_{p}^{l-1}\;=\;H_{1}|U_{p}^{l-1}$ $\;=\;p^{(2s-1)(l-1)}H_{l}$
$\;\equiv\;p^{(2s-1)(l-1)}\big(G_{l}|T_{p,2s}-G_{l-1}\big)$
$\;=\;p^{(2s-1)l}G_{l+1}\;\equiv\;0\pmod{p^{(2s-1)l}}$.
@@
Let $2s\in\{4,6,8,10,14\}$ and let $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$. Let $m\in\mathbb{Z}$ with $m\geq 0$ and $(-1)^{s-1}m\equiv 0,1\pmod{4}$ and let $p$ be a prime. Let
$F_{m}\;:=\;\mathcal{S}_{d_{0}}f_{s+1/2,m}.$
Then, for all $n,l\in\mathbb{Z}^{+}$,
$p^{(s-1)t}\cdot a_{np^{l}}(F_{m})\;\equiv\;0\pmod{p^{(s-1)l}}$,
where $t$ is the largest integer such that $p^{2t}{\hskip 1.0pt}|{\hskip 1.0pt}m$ and $(-1)^{s-1}{\hskip 1.0pt}m/p^{2t}\equiv 0,1\pmod{4}$.
@@
We will treat $t\geq 0$ as a variable and prove this by induction on $t$. Note that the base case when $t=0$ follows from Proposition 6.5. Let $m^{\prime}=m/p^{2t}$. Then, $(-1)^{s-1}m^{\prime}\equiv 0,1\pmod{4}$ and the pair $(p,m^{\prime})$ satisfies the assumption in Proposition 6.5. Define $G_{i}$ as in the proof of Proposition 6.5 with respect to $m^{\prime}$. Then, by Lemma 6.4,
$G_{i+1}$ $\;=\;\mathcal{S}_{d_{0}}f_{s+1/2,{\hskip 1.0pt}m^{\prime}p^{2i}}\;=\;F_{m^{\prime}p^{2i}}.$
Let $H_{i}=G_{i}|U_{p}-G_{i-1}$ as before. Then,
$G_{t+1}|U_{p}^{l}\;=\;(H_{t+1}+G_{t})|U_{p}^{l-1}\;=\;p^{-(2s-1)t}H_{1}|U_{p}^{t+l-1}+G_{t}|U_{p}^{l-1}.$
Suppose that $t\geq 1$. Note that by the last equation in the proof of Proposition 6.5,
$H_{1}|U_{p}^{t+l-1}\;\equiv\;0\pmod{p^{(2s-1)(t+l)}}$
and that by induction hypothesis,
$p^{(s-1)(t-1)}\cdot G_{t}|U_{p}^{l-1}\;\equiv\;0\pmod{p^{(s-1)(l-1)}}$.
It follows that
$p^{(s-1)t}\cdot G_{t+1}|U_{p}^{l}\;\equiv\;0\pmod{p^{(s-1)l}}$.
The result then follows. ∎
@@
Let $2s\in\{4,6,8,10,14\}$ and let $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$. Let $m\in\mathbb{Z}$ with $m\geq 0$ and $(-1)^{s-1}m\equiv 0,1\pmod{4}$ and let $A$ be the largest integer such that $A^{2}{\hskip 1.0pt}|{\hskip 1.0pt}m$ and $(-1)^{s-1}m/A^{2}\equiv 0,1\pmod{4}$. Then,
@@
is (s1)(s-1)-magnetic, i.e., $n^{s-1}{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}(F)$ for all $n\in\mathbb{Z}^{+}$.

### v2

@@ 7.3. Hypergeometric reinterpretation of $a_{p^{l}}(F_{4,C})$
Now, one can relate $a_{\mathfrak{p}}(C)$ and $a_{\mathfrak{P}}(C_{1})$ via the fact that $j(C)=j(C_{1})$ and use an appropriate version of the truncated Clausen’s formula to obtain the desired congruence.
8. Proof of Theorem 1.10
We will now prove Theorem 1.10, i.e. the supercongruences and the $\frac{k-2}{2}$-magnetic property of $G_{k,D}^{(k/2)}$, or rather its generalization Theorem 5.10. In fact, we will further generalize this to a result that holds for any discriminant $D<0$.
Recall that for $k\in 2\mathbb{Z}_{\geq 2}$,
$G_{k}(z,\tau)\;=\;\frac{g_{k}(z)\cdot g_{2-k}(\tau)}{j(z)-j(\tau)}.$
For a discriminant $D<0$, define
$\mathcal{G}_{k,D}(z)\;:=\;\frac{2}{w_{D}}\sum_{Q\in\mathcal{Q}_{D}^{\mathrm{prim}}}(\partial^{\frac{k-2}{2}}_{\tau}G_{k})(z,\alpha_{Q}).$ (27)
As before, since $(\partial^{\frac{k-2}{2}}_{\tau}G_{k})(z,\tau)$ is of weight 0 with respect to $\tau$, $\mathcal{G}_{k,D}(z)$ is independent of the choice of $\mathcal{Q}_{D}^{\mathrm{prim}}$.
@@
Let $k\in 2\mathbb{Z}_{\geq 2}$ and $\lambda$ be a relation for $S_{k}$. Let $D<0$ be a discriminant and write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}$ and some fundamental discriminant $D_{0}<0$. Let
[…]
Then, $\widetilde{\mathcal{G}}_{k,D}$ satisfies the following.
@@
    $\widetilde{\mathcal{G}}_{k,D}\in\mathbb{Z}[[q]]$ and $\widetilde{\mathcal{G}}_{k,D}|\lambda$ is $\frac{k-2}{2}$-magnetic, i.e., for all $n\in\mathbb{Z}^{+}$,
    $n^{\frac{k-2}{2}}{\hskip 1.0pt}\big|{\hskip 1.0pt}a_{n}(\widetilde{\mathcal{G}}_{k,D}|\lambda).$
@@
    $a_{np^{l}}(\widetilde{\mathcal{G}}_{k,D}|\lambda)\;\equiv\;\big(\big(\tfrac{D}{p}\big)p\big)^{\frac{k-2}{2}}a_{np^{l-1}}(\widetilde{\mathcal{G}}_{k,D}|\lambda)\pmod{p^{(k-1)l}}$.
Remark 8.2.
This theorem should be viewed as a natural generalization of Li–Neururer [LN19, Theorem 1.5] and Paşol–Zudilin [PZ22, Theorems 1 and 2]. Indeed, the proof uses the same method: one considers the half-integral weight preimage of $\mathcal{G}_{k,D}$ under the Shimura lift, analyzes the $U_{p}$ action on the preimage, and transfers the properties of the preimage to the supercongruences and the magnetic property of $\mathcal{G}_{k,D}$.
@@ 8.1. The Shimura lift
Let $s\in\mathbb{Z}$, $f\in M_{s+1/2}^{!,+}$, and $d_{0}$ be a fundamental discriminant $d_{0}$ with $(-1)^{s}d_{0}>0$. Define the $d_{0}$-th Shimura lift of $f$ by
@@
Now, let $d$ be a discriminant and $d_{0}$ be a fundamental discriminant with $dd_{0}<0$. For an $\mathrm{SL}_{2}(\mathbb{Z})$-invariant function $f(\tau)$ on $\mathcal{H}$, recall the following definition from [DJ08b]
$\mathrm{Tr}_{d,d_{0}}(f)\;:=\;\sum_{Q\in\mathcal{Q}_{dd_{0}}}\frac{1}{w_{Q}}\cdot\chi(Q)f(\alpha_{Q}),$ (28)
where for each $Q(X,Y)=aX^{2}+bXY+cY^{2}\in\mathcal{Q}_{dd_{0}}$, we define
@@
Now, let $s\in\mathbb{Z}_{\geq 2}$, $d$ be a discriminant, and $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$ and $dd_{0}<0$. Write $2s=k^{\prime}+12\delta$ for some $k^{\prime}\in\{0,4,6,8,10,14\}$ and $\delta\in\mathbb{Z}$. We define
$h_{s,d,d_{0}}\;:=\;\sum_{n=0}^{\delta}a_{n}\big(\mathcal{S}_{d_{0}}f_{s+1/2,|d|}\big)\cdot g_{2s,-n}.$
Lemma 8.3.
Let $s\in\mathbb{Z}_{\geq 2}$, $d$ be a discriminant, and $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$ and $dd_{0}<0$. Then, $h_{s,d,d_{0}}$ has integer coefficients and lies in $S_{2s}$.
Proof.
The integrality of $h_{s,d,d_{0}}$ simply follows from the integrality of $f_{s+1/2,|d|}$ and $g_{2s,-n}$. Now, since $s\in\mathbb{Z}_{\geq 2}$ and $|d|>0$, it is easy to check that $a_{0}(f_{s+1/2,|d|})=0$, so $a_{0}\big(\mathcal{S}_{d_{0}}f_{s+1/2,|d|}\big)=0$. Hence, $h_{s,d,d_{0}}\in S_{2s}$. ∎
@@
Let $s\in\mathbb{Z}_{\geq 2}$, $d$ be a discriminant, and $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$ and $dd_{0}<0$. Then,
$\big(\mathcal{S}_{d_{0}}f_{s+1/2,|d|}\big)(z)-h_{s,d,d_{0}}(z)\;=\;-(-1)^{\lfloor\frac{s-1}{2}\rfloor}{\hskip 1.0pt}|d|^{-\frac{s}{2}}{\hskip 1.0pt}|d_{0}|^{\frac{s-1}{2}}\;\mathrm{Tr}_{d,d_{0}}\big((\partial^{s-1}_{\tau}G_{2s})(z,\tau)\big),$
@@
See [DJ08b, Proposition 10]. ∎
8.2. The Hecke actions
We will now analyze the actions of the Hecke operators on $f_{s+1/2,m}$ and transfer them to modular forms of integral weight.
@@ 8.2. The Hecke actions
Let $s\in\mathbb{Z}_{\geq 2}$, $m\in\mathbb{Z}^{+}$ with $(-1)^{s-1}m\equiv 0,1\pmod{4}$, and $p$ be a prime such that
@@
$g_{0}$ $\;:=\;f_{s+1/2,m}$
$g_{1}$ $\;:=\;g_{0}|T_{p,s+1/2}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}\cdot g_{0}$
$g_{i+1}$ $\;:=\;g_{i}|T_{p,s+1/2}-p^{2s-1}\cdot g_{i-1}$
@@
$g_{i}\;=\;p^{(2s-1)i}q^{-mp^{2i}}+O(q)\in\mathbb{Z}[[q,q^{-1}]].$
In particular, $g_{i}-p^{(2s-1)i}\cdot f_{s+1/2,mp^{2i}}$ has integer coefficients and lies in $S_{s+1/2}^{+}$.
@@
Write $f_{m}=f_{s+1/2,m}$. The case when $i=0$ follows from the definition of $f_{m}$. For $i=1$, it is easy to check that the assumptions on $p$ and $m$ imply that
$g_{1}$ $\;=\;g_{0}|T_{p,s+1/2}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}\cdot g_{0}$
$\;=\;\left(p^{2s-1}q^{-mp^{2}}+\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}q^{-m}+O(q)\right)-\left(\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}q^{-m}+O(q)\right)$
$\;=\;p^{2s-1}q^{-mp^{2}}+O(q),$
from which the result follows. Now, for $i\geq 1$, it is easy to check that by induction hypothesis,
$g_{i+1}\;=\;g_{i}|T_{p,s+1/2}-p^{2s-1}\cdot g_{i-1}\;=\;p^{(2s-1)(i+1)}q^{-mp^{2i+2}}+O(q).$
The result then follows. ∎
Now, recall that the Shimura lift is Hecke-equivariant. That is, for $s\in\mathbb{Z}$, $f\in M_{s+1/2}^{!,+}$, and a rational prime $p$
$\mathcal{S}_{d_{0}}(f|T_{p,s+1/2})\;=\;\mathcal{S}_{d_{0}}(f)|T_{p,2s},$
@@
Let $s\in\mathbb{Z}_{\geq 2}$ and let $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$. Let $m\in\mathbb{Z}^{+}$ with $(-1)^{s-1}m\equiv 0,1\pmod{4}$ and let $p$ be a prime such that
@@
and let $\lambda$ be a relation for $S_{2s}$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{m}|\lambda)\;\equiv\;\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}\cdot a_{np^{l-1}}(F_{m}|\lambda)\pmod{p^{(2s-1)l}}$.
@@
$G_{0}$ $\;=\;F_{m}$
$G_{1}$ $\;=\;G_{0}|T_{p,2s}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}\cdot G_{0}$
$G_{i+1}$ $\;=\;G_{i}|T_{p,2s}-p^{2s-1}\cdot G_{i-1}.$
Let $\widetilde{G}_{i}=G_{i}|\lambda$. Then, since $\lambda$ commutes with $T_{p,2s}$, we have
$\widetilde{G}_{i+1}$ $\;=\;\widetilde{G}_{i}|T_{p,2s}-p^{2s-1}\cdot\widetilde{G}_{i-1}.$
Write $H_{i}=\widetilde{G}_{i}|U_{p}-p^{2s-1}\widetilde{G}_{i-1}$. Then,
$H_{i+1}\;=\;\widetilde{G}_{i+1}|U_{p}-p^{2s-1}\widetilde{G}_{i}\;=\;\left(\widetilde{G}_{i}|(U_{p}+p^{2s-1}V_{p})-p^{2s-1}\widetilde{G}_{i-1}\right)\hskip-2.27626pt\big|U_{p}-p^{2s-1}\widetilde{G}_{i}\;=\;H_{i}|U_{p}.$
It follows that
$H_{1}|U_{p}^{l-1}\;=\;H_{l}.$
Now, since the Shimura lift preserves cusp forms and $\lambda$ annihilates cusp forms, it follows from Lemma 8.5 that
$\widetilde{G}_{i}\;=\;p^{(2s-1)i}(\mathcal{S}f_{s+1/2,mp^{2i}})|\lambda\;\equiv\;0\pmod{p^{(2s-1)i}}$.
@@
$H_{l}\;=\;\widetilde{G}_{l}|U_{p}-p^{2s-1}\widetilde{G}_{l-1}\;\equiv\;\widetilde{G}_{l}|T_{p,2s}-p^{2s-1}\widetilde{G}_{l-1}\;\equiv\;\widetilde{G}_{l+1}\;\equiv\;0\pmod{p^{(2s-1)(l+1)}}$. (29)
Now,
$F_{m}|\lambda U_{p}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}\cdot F_{m}|\lambda$ $\;=\;\widetilde{G}_{0}|U_{p}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}G_{0}$
$\;=\;\widetilde{G}_{1}-p^{2s-1}\widetilde{G}_{0}|V_{p}\;\equiv\;0\pmod{p^{2s-1}}$,
and for $l\geq 2$,
$F_{m}|\lambda U_{p}^{l}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}\cdot F_{m}|\lambda U_{p}^{l-1}$ $\;=\;\widetilde{G}_{1}|U_{p}^{l-1}-p^{2s-1}\widetilde{G}_{0}|U_{p}^{l-2}$
$\;=\;H_{1}|U_{p}^{l-2}$
$\;=\;H_{l-1}\;\equiv\;0\pmod{p^{(2s-1)(l+1)}}$.
@@
Let $s\in\mathbb{Z}_{\geq 2}$ and let $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$. Let $m\in\mathbb{Z}^{+}$ with $(-1)^{s-1}m\equiv 0,1\pmod{4}$ and let $p$ be a prime such that
  1. (1)
    $p^{2}\nmid m$, or
  2. (2)
Let $t\in\mathbb{Z}$ with $t\geq 0$ and let
$F_{mp^{2t}}\;:=\;\mathcal{S}_{d_{0}}f_{s+1/2,mp^{2t}}.$
Let $\lambda$ be a relation for $S_{2s}$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{mp^{2t}}|\lambda)\;\equiv\;\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}\cdot a_{np^{l-1}}(F_{mp^{2t}}|\lambda)\pmod{p^{(2s-1)(l-t)}}$.
In particular, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{mp^{2t}}|\lambda)\;\equiv\;0\pmod{p^{(s-1)(l-t)}}$.
@@
We will prove this by induction on $t$. Note that the base case when $t=0$ follows from Proposition 8.6. $G_{i}$ and $\widetilde{G}_{i}=G_{i}|\lambda$ as in the proof of Proposition 8.6 with respect to $m$. Since the Shimura lift preserves cusp forms and $\lambda$ annihilates cusp forms, it follows from Lemma 8.5 that
$\widetilde{G}_{i}\;=\;p^{(2s-1)i}(\mathcal{S}_{d_{0}}f_{s+1/2,mp^{2i}})|\lambda\;=\;p^{(2s-1)i}F_{mp^{2i}}|\lambda.$
Let $H_{i}=\widetilde{G}_{i}|U_{p}-p^{2s-1}\widetilde{G}_{i-1}$ as before. Then,
$\widetilde{G}_{t}|U_{p}^{l}\;=\;(H_{t}+p^{2s-1}\widetilde{G}_{t-1})|U_{p}^{l-1}\;=\;H_{t+l-1}+p^{2s-1}\widetilde{G}_{t-1}|U_{p}^{l-1}.$
By (29),
$H_{t+l-1}\;\equiv\;0\pmod{p^{(2s-1)(t+l)}}$.
Hence,
$F_{mp^{2t}}|\lambda U_{p}^{l}\;=\;p^{-(2s-1)t}H_{t+l-1}+F_{mp^{2t-2}}|\lambda U_{p}^{l-1}\;\equiv\;F_{mp^{2t-2}}|\lambda U_{p}^{l-1}\pmod{p^{(2s-1)l}}$,
so
$F_{mp^{2t}}|\lambda U_{p}^{l}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}F_{mp^{2t}}|\lambda U_{p}^{l-1}\;\equiv\;$
$\hskip 56.9055ptF_{mp^{2t-2}}|\lambda U_{p}^{l-1}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}F_{mp^{2t-2}}|\lambda U_{p}^{l-2}\pmod{p^{(2s-1)l}}$.
The result now follows from induction. ∎
@@
Let $s\in\mathbb{Z}_{\geq 2}$ and let $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$. Let $m\in\mathbb{Z}^{+}$ with $(-1)^{s-1}m\equiv 0,1\pmod{4}$ and let $A$ be the largest integer such that $A^{2}{\hskip 1.0pt}|{\hskip 1.0pt}m$ and $(-1)^{s-1}m/A^{2}\equiv 0,1\pmod{4}$. Let
@@
and let $\lambda$ be a relation for $S_{2s}$. Then, $F|\lambda$ is (s1)(s-1)-magnetic, i.e., $n^{s-1}{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}(F|\lambda)$ for all $n\in\mathbb{Z}^{+}$.